OpenStudy (anonymous):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term 3? A. An = 5-3(n-1); all integers where n≥1 B. An = 5-3(n-1); all integers where n≥0 C. An = 5-2(n-1); all integers where n≥0 D. An = 5-2(n-1); all integers where n≥1

OpenStudy (anonymous):

@omarbirjas do you know what the answer is or do you know how to solve it?

OpenStudy (xapproachesinfinity):

any arithmetic sequence can be written as \[A_n=A+(n-1)d\] where A is the first term and d is common difference

OpenStudy (xapproachesinfinity):

to find d all you do is next term-previous term

OpenStudy (xapproachesinfinity):

in this case 3-5=-2 d=-2

OpenStudy (xapproachesinfinity):

first term you have it it is 5

OpenStudy (xapproachesinfinity):

now do the rest, in fact i did give everything lol

OpenStudy (anonymous):

so it would be An= 5 +(n-1) -2

OpenStudy (xapproachesinfinity):

yes

OpenStudy (anonymous):

I think the is answer A? is that right?

OpenStudy (xapproachesinfinity):

no

OpenStudy (anonymous):

darn it

OpenStudy (xapproachesinfinity):

that does not match A

OpenStudy (anonymous):

wait no it has to be D because of the -2

OpenStudy (xapproachesinfinity):

well the way you wrote and the way d is written are the same they just moved -2 from instead C could be the answer according to your line of reasoning because c has -2 as well

OpenStudy (xapproachesinfinity):

so what determines the correct answer is that condition tied to it n>=1 for d and n>=0 for c which is correct for our case for the first term to be 5

OpenStudy (anonymous):

so the n≥1 is wrong

OpenStudy (xapproachesinfinity):

it is d because if it is c the first term would have been 7 not 5

OpenStudy (xapproachesinfinity):

it is d because when you play n=1 for the first term what you get is A1=5-2(1-1)=5-0=5

OpenStudy (xapproachesinfinity):

plug* i meant not play

OpenStudy (anonymous):

I'm sorry I was confused! Thank you so much! This helped me a lot!

OpenStudy (xapproachesinfinity):

no problem